OpenStudy (anonymous):

derivative of y= (x^4+3x^2-2)^5

OpenStudy (cwrw238):

you us ethe chain rule - are you familiar with this?

OpenStudy (anonymous):

f'(u^n) = nu'u(n-1) F'(x) = 5 (4x^3 + 6x) (x^4 + 3x^2 -2)^4

OpenStudy (cwrw238):

[f(g(x)]' = f'g(x) * g'(x)

OpenStudy (anonymous):

but the answer in the bookis diff

OpenStudy (anonymous):

10 x (3+2 x^2) (-2+3 x^2+x^4)^4

OpenStudy (anonymous):

i got your asnwers, but i dont undestnad how that becomes what the book says

OpenStudy (anonymous):

5(4x^3+6x)^4 (20x^3+30x)^4 (5*(x^4+3x^2-2)^4)*(4x^3+6x)

OpenStudy (cwrw238):

jordan's answer is correct but it can be simplidied 10x(2x^2 + 3)(x^4 + 3x^2 - 2)^4

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

sorry about that

OpenStudy (anonymous):

@cwrw238 how do u simplify it. can you do it step by step. i dont see how it becomes that final answer

OpenStudy (cwrw238):

the first part can be simplified 5(4x^3 + 6x) take the part in the parentheses the greatest common factor is 2x right?

OpenStudy (anonymous):

oh. ya, okay.

OpenStudy (cwrw238):

so we can write it as 2x(2x^2 + 3)

OpenStudy (cwrw238):

this multiplied by the 5 gives 10x(2x^2 + 3)

OpenStudy (anonymous):

thank you so much for your help!

OpenStudy (cwrw238):

yw